Solve system of equation by elimination:
5b = 20 + 2a
2a + 4b = 7

Found 3 solutions by Boreal, Alan3354, Assumpcaod:
Answer by Boreal(15235)

5b = 20 + 2a
2a + 4b = 7
move the a to the other side with b
-2a+5b=20
2a+4b=7
If you add them, the a will disappear. It isn't always this convenient, but here it is.
9b=27
b=3
substitute in the first equation and 15=20+2a
a=-2 1/2
substitute into the second to check.
-5+12=7
that works
a is -2 1/2
b is 3

Answer by Alan3354(69443)

5b = 20 + 2a --> 2a - 5b = -20
2a + 4b = 7
-------------
2a - 5b = -20
2a + 4b = 7
--------------------- Subtract

Answer by Assumpcaod(1)